OpenStudy (ledah):

The total number of degrees in a regular hexagon is 720°. If a corresponding angle in a congruent regular hexagon is (11x – 12)°, find the value of x.

ganeshie8 (ganeshie8):

a regular hexagon has 6 congruent sides, 6 congruent interior angles. so we can use this, interior angle = 720/6 = (11x-12) => 120 = 11x-12 x = ?

OpenStudy (ledah):

add 12 to each side?

OpenStudy (ledah):

x=12 is that right?